# Cross-Lingual Multi-Hop Knowledge Editing – Benchmarks, Analysis and a Simple Contrastive Learning based Approach

### v1

@@ Abstract
Large language models are often expected to constantly adapt to new sources of knowledge and knowledge editing techniques aim to efficiently patch the outdated model knowledge, with minimal modification. Most prior works focus on monolingual knowledge editing in English, even though new information can emerge in any language from any part of the world. We propose the Cross-Lingual Multi-Hop Knowledge Editing paradigm, for measuring and analyzing the performance of various SoTA knowledge editing techniques in a cross-lingual setup. Specifically, we create a parallel cross-lingual benchmark, CroLin-MQuAKE for measuring the knowledge editing capabilities. Our extensive analysis over various knowledge editing techniques uncover significant gaps in performance between the cross-lingual and English-centric setting. Following this, we propose a significantly improved system for cross-lingual multi-hop knowledge editing, CLeVer-CKE. CLeVer-CKE is based on a retrieve, verify and generate knowledge editing framework, where a retriever is formulated to recall edited facts and support an LLM to adhere to knowledge edits. We develop language-aware and hard-negative based contrastive objectives for improving the cross-lingual and fine-grained fact retrieval and verification process used in this framework. Extensive experiments on three LLMs, eight languages, and two datasets show CLeVer-CKE’s significant gains of up to 30% over prior methods.
@@ 1 Introduction
Prior work on knowledge editing has largely focused on editing LLMs in a monolingual setting (Zhong et al., 2023; Gu et al., 2024), where both user queries and edited facts are expressed in the form of English. These works can be grouped into two categories: parameter-update and parameter-preserving methods. The former directly updates the parameters within LLMs for updating knowledge about the edited facts through meta-learning, fine-tuning, or knowledge locating (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022a; Meng et al., 2022a, b). The later approach freezes the parameters and explicitly stores the edited facts in an external memory and retrieves them for answering user queries (Zhong et al., 2023; Gu et al., 2024; Mitchell et al., 2022c; Hartvigsen et al., 2023). Existing monolingual knowledge editing techniques aren’t broadly applicable since new knowledge can emerge in different languages. Some works have made progress in this direction (Beniwal et al., 2024; Xu et al., 2023a; Si et al., 2024), but they have considered a simplistic setting of assuming the edited facts as independent without any multi-hop rippling consequences on entailed reasoning process, and are primarily focused on parameter-modifying based editing methods.
There has only been a limited focus on the realistic case of cross-lingual multi-hop knowledge editing (see Fig 1), where the edited knowledge can come in through users who communicate in different languages. Further, much of edited knowledge often has a rippling effect on other facts of the world. For example, the club change of Messi affects deduction process of question “indicating a superficial word matching rather than a contextual grasp of the entities involved." This knowledge editing setting, which we argue is important to study, is challenging since the model needs to transfer knowledge about fact edits between different languages, while also reasoning about the facts which are modified as a consequence to the given edit. Poor knowledge transfer between languages can lead to error propagation across reasoning steps which can increase failure cases of model editing.
@@ 2 Cross-lingual Multi-hop Editing
For evaluating models in a cross-lingual multi-hop setting, we make use of multi-hop questions such as $Q_{L_{n}}$, in language $L_{n}$ which is different from $L_{1\dots k}$. The question asks about the head entity $s_{1}^{L_{1}}$ for which the answer is $o_{n}^{L_{k}}$ before editing. After editing, the fact chain changes to $\mathcal{P^{*}}=\langle(s_{1}^{L_{1}},r_{1}^{L_{1}},o_{1}^{{L_{1}}*})$ $,(s_{2}^{L_{2}},r_{2}^{L_{2}},o_{2}^{{L_{2}}*}),\dots,(s_{n}^{L_{k}},r_{n}^{L_{k}},o_{n}^{{L_{k}}*})\rangle$ since edits in the first fact can effect the subsequent facts it’s linked to. For answering $Q_{L_{n}}$ after editing, the model has to account for this rippling effect, and provide the final answer as $o_{n}^{{L_{k}}*}$. For this, model has to transfer knowledge of the edited fact and the answer, between languages $L_{1\dots k}$ and $L_{n}$, while correctly reasoning about fact edits via $\mathcal{P^{*}}$.
@@ 3 CroLin-MQuAKE Benchmark
We develop one of the first parallel cross-lingual for measuring the knowledge editing capabilities of the existing approaches. A parallel benchmark has the same test examples across all the languages, enabling a direct comparison between them. For this, we use existing datasets measuring the multi-hop model editing in English: MQuAKE-CF and MQuAKE-T released by Zhong et al. (2023), which have counterfactual edits and real-world temporal edits respectively. We translate one fact edit per example in these datasets using Google Translate (Google, ) into 7 languages with diverse writing scripts across medium to high resourcedness - German, Spanish, Chinese, Rissian, Hindi, Bengali, Swahili. This results in the benchmark: Cross-Lingual Multi-Hop QnA for Knowledge Editing (CroLin-MQuAKE). It has two datasets, CroLin-MQuAKE-CF and CroLin-MQuAKE-T, each having 8 languages, and 3k and 1.8k parallel examples (same examples in all languages) per language, respectively. The translations are verified by human experts proficient in particular languages and evaluation of BLEU score Papineni et al. (2002) using backtranslation. We find that the translation is highly accurate, since we study medium to high resource languages. See Section A.2 for more details.
@@ 4 Benchmark Analysis on Cross-Lingual Multi-hop Knowledge Editing
Evaluation Metrics. We use multi-hop accuracy proposed by Zhong et al. (2023) which measures the accuracy of the final answer of a multi-hop question. We also adopt hop-wise answering accuracy for checking the correctness of intermediate reasoning steps, as proposed by Gu et al. (2024).
@@
  • PokeMQA-CL: PokeMQA is similar to MeLLo but consists of a conflict disambiguator for retrieving as well as classifying if a fact is useful to answer a sub-question. Following PokeMQA, we train this disambiguator using BCE loss with negative sampling for retrieving the close edits, given a decomposed sub-question. However, our training dataset now consists of translated version of the training dataset used in PokeMQA. This training set contains all 8 languages (the multilingual setting) or English along with one of the 7 non-English languages (the bilingual setting).
@@
Knowledge editing performance based on retriever training technique. MeLLo-CL retrieves the edited fact from the memory using mContriever and employs an LLM to disambiguate between the generated answer and the retrieved fact and hence ascertains if the generated fact needs any update or not. On the other hand, the current state-of-the-art knowledge editing method in English, PokeMQA-CL, uses a retrieve-then-verify approach, which offloads the knowledge disambiguation to the retriever. This retriever is a light-weight and fine-tuned distilbert-base model trained on a (sub-question,edit) pair dataset using binary cross-entropy loss with negative sampling. It retrieves the closest edits (in fact memory) to a sub-question and scores it for whether the edit answers the question or not (called verification or disambiguation). If it does, then it uses this new knowledge as the answer to the sub-question in the n-th hop step and performs in-context editing. PokeMQA-CL outperforms MeLLo-CL on in the monolingual (English) setting, with a much smaller retriever as shown in Gu et al. (2024), however, when trained with multilingual data, we find that it significantly under-performs MeLLo-CL in most languages including English as shown in Fig. 2. MeLLo-CL under-performs in Hindi and Bengali – languages with scripts very different from Latin, even though it’s retriever is trained with 100+ languages.
@@
❶ MeLLo-CL: When edits are made in English, 63.3% of the samples are correct, 29.3% have the contradiction error, 3.6% have Incorrect retrieval, and 3.6% have the incorrect LLM response. For edits made in Hindi, 33.3% of the samples are correct, 60% exhibit an error combination of incorrect retrieval and subsequent contradiction error, where the model first makes an incorrect retrieval and then fails in the contradiction step and 6.6% of erroneous samples are due to the incorrect LLM response. In the CroLin-MQuAKE-CF case when the multilingual edited fact memory containing edits in English and Hindi, MeLLo-CL’s retriever rarely retrieves edits in Hindi, indicating a limitation in its multilingual capabilities. The limitation of MeLLo-CL lies in its retriever-then-contradict mechanism which is up to the LLM.
❷ PokeMQA-CL: When edits are made in English, 53.3% of the samples are correct and 46.3% have the incorrect retrieval error. When edits are made in Hindi, 43.3% are correct, 51% have errors due to the incorrect retrieval and 5.6% are due to the incorrect LLM response. The limitation of PokeMQA-CL lies in its reliance on a bag-of-words model for retrieval. For instance, when presented with the sub-question “Who is the head of state of the USA?", it retrieves the fact “The head of state of Mongolia is Khürelsükh Ukhnaa." This example underscores that PokeMQA-CL prioritizes facts with the highest word overlap, specifically “head of state" indicating a superficial word matching rather than a contextual grasp of the entities involved.
@@ 5 CLeVer-CKE for Knowledge Editing
For overcoming limitations in cross-lingual multi-hop knowledge editing, we design CLeVer-CKE, a cross-lingual and light-weight model editor that seamlessly integrates into any backbone LLM, without changing its parameters. CLeVer-CKE is inspired by memory-based and retrieval-augmented knowledge editing methods (Zhong et al., 2023; Gu et al., 2024; Mitchell et al., 2022b) for mutlihop question answering. CLeVer-CKE follows the following procedure: Given an input query, it a) decomposes the multi-hop question into multiple sub-questions for getting to the final answer, and for answering each sub-question b) retrieves a relevant fact from the edit memory, c) disambiguates whether the retrieved new knowledge is relevant to answering the sub-question, and d) continues the model generation process based on that. In this work, we primarily aim at showing the importance of having a high-quality retriever for the retrieve-and-verify steps at b) and c) described as follows. See Fig. 3 for an overview.
@@
BCE Loss: Following (Gu et al., 2024; Mikolov et al., 2013) we add a binary cross-entropy loss in the cross-lingual setting as a baseline loss for training the retriever for retrieving edits in a cross-lingual setting. The negative BCE Loss function takes questions in English and their corresponding edited facts in one of the seven languages as input. We then compute the $L_{2}$ norm between these edits and questions, and sample 20 negatives. The loss function $\mathcal{L}$ is defined similar to Gu et al. (2024):
@@ 6 Related Works
Cross-lingual knowledge editing. Recent studies have shifted focus to the multilingual capabilities of SoTA LLMs like LLaMA Touvron et al. (2023a), ChatGPT Schulman et al. (2022), and GPT-4 OpenAI (2023). Wang et al. (2023a) investigated cross-lingual knowledge editing and its impact on different target languages using a synthetic dataset. Si et al. (2024) introduced Multilingual Patch Neuron (MPN) for efficient cross-lingual knowledge synchronization, showing enhanced performance on single-hop XNLI and XFEVER datasets. Xu et al. (2023b) proposed a framework for language anisotropic editing, facilitating simultaneous cross-lingual model editing. Beniwal et al. (2024) explored the cross-lingual model editing (XME) paradigm, revealing performance limitations in multilingual LLMs for hypernetwrok based parameter-modifying methods. Wang et al. (2023b) presented Retrieval-augmented Multilingual Knowledge Editing (ReMaKE), a model-agnostic knowledge editing method designed for multilingual settings. ReMaKE retrieves new knowledge from a multilingual knowledge base and concatenates it with prompts to update LLMs. Most works assume edited facts are independent without any multi-hop consequences of these edits, and focus on parameter update based methods. We focus on parameter-preserving methods, and the more complex setting of multi-hop editing in a cross-lingual setup. See A.1 for more.
@@ Appendix A Appendix
A.1 Related Work
Knowledge editing methods: Knowledge editing can be broadly classified intro two groups. 1) Parameter-modifying based editing which locates the parameters related to factual knowledge and subsequently modify them (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022a; Meng et al., 2022a, b). These method requires an error-prone analytic step to identify parameters, which might be model-specific and not efficient. 2) Parameter-preserving based editing keeps the model parameters frozen and explicitly stores the fact edits in an external memory, for retrieval and external validation (Zhong et al., 2023; Gu et al., 2024; Mitchell et al., 2022c; Hartvigsen et al., 2023). some recent works like that of Hernandez et al. (2023) have also explored a decoding time approach for editing knowledge.
A.2 Verification of Translated Data in CroLin-MQuAKE
A.2.1 Human Verification of Translation
@@
A.2.2 Verification of Translations via Backtranslation
@@
A.3 Training Details
We employ the training dataset to train the retriever component of the CLeVer-CKE framework, using the same training set as utilized in training PokeMQA-CL Gu et al. (2024). Subsequently, we translate this dataset into seven other languages and generate hard negatives following the method outlined in Section 5. The training dataset contains 6688 samples along with translations into 8 langugaes and hard-negative pairs for each edit in the dataset, both of which is created by us for training CLeVer-CKE’s retriever. For training the multilingual retriever, we utilize data from all languages, while for training the bilingual retriever, we focus on English and the target language data. To optimize our method’s performance, we conduct hyperparameter tuning on a validation set derived from CroLin-MQuAKE-CF, comprising 50 samples exclusively for this purpose without involvement in inferencing tasks. The hyperparameters used for tuning are mentioned in Table 4. Our experiments are expensive (See Appendix  A.7) and we do not perform experiments on multiple seeds.
A.4 Method Details

### v2

@@ Abstract
Large language models are often expected to constantly adapt to new sources of knowledge and knowledge editing techniques aim to efficiently patch the outdated model knowledge, with minimal modification. Most prior works focus on monolingual knowledge editing in English, even though new information can emerge in any language from any part of the world. We propose the Cross-Lingual Multi-Hop Knowledge Editing paradigm, for measuring and analyzing the performance of various SoTA knowledge editing techniques in a cross-lingual setup. Specifically, we create a parallel cross-lingual benchmark, CroLin-MQuAKE for measuring the knowledge editing capabilities. Our extensive analysis over various knowledge editing techniques uncover significant gaps in performance between the cross-lingual and English-centric setting. Following this, we propose a significantly improved system for cross-lingual multi-hop knowledge editing, CLeVer-CKE. CLeVer-CKE is based on a retrieve, verify and generate knowledge editing framework, where a retriever is formulated to recall edited facts and support an LLM to adhere to knowledge edits. We develop language-aware and hard-negative based contrastive objectives for improving the cross-lingual and fine-grained fact retrieval and verification process used in this framework. Extensive experiments on three LLMs, eight languages, and two datasets show CLeVer-CKE’s significant gains of up to 30% over prior methods. Code and data are released at https://github.com/HarmanDotpy/CroLin-KE
@@ 1 Introduction
Prior work on knowledge editing has largely focused on editing LLMs in a monolingual setting (Zhong et al., 2023; Gu et al., 2024a), where both user queries and edited facts are expressed in the form of English. These works can be grouped into two categories: parameter-update and parameter-preserving methods. The former directly updates the parameters within LLMs for updating knowledge about the edited facts through meta-learning, fine-tuning, or knowledge locating (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022a; Meng et al., 2022a, b). The later approach freezes the parameters and explicitly stores the edited facts in an external memory and retrieves them for answering user queries (Zhong et al., 2023; Gu et al., 2024a; Mitchell et al., 2022c; Hartvigsen et al., 2023). Existing monolingual knowledge editing techniques aren’t broadly applicable since new knowledge can emerge in different languages. Some works have made progress in this direction (Beniwal et al., 2024; Xu et al., 2023a; Si et al., 2024), but they have considered a simplistic setting of assuming the edited facts as independent without any multi-hop rippling consequences on entailed reasoning process, and are primarily focused on parameter-modifying based editing methods.
There has only been a limited focus on the realistic case of cross-lingual multi-hop knowledge editing (see Fig 1), where the edited knowledge can come in through users who communicate in different languages. Further, much of edited knowledge often has a rippling effect on other facts of the world. For example, the club change of Messi affects deduction process of question "indicating a superficial word matching rather than a contextual grasp of the entities involved." This knowledge editing setting, which we argue is important to study, is challenging since the model needs to transfer knowledge about fact edits between different languages, while also reasoning about the facts which are modified as a consequence to the given edit. Poor knowledge transfer between languages can lead to error propagation across reasoning steps which can increase failure cases of model editing.
@@ 2 Cross-lingual Multi-hop Editing
For evaluating models in a cross-lingual multi-hop setting, we make use of multi-hop questions such as $Q_{L_{n}}$, in language $L_{n}$ which is different from $L_{1\dots k}$. The question asks about the head entity $s_{1}^{L_{1}}$ for which the answer is $o_{n}^{L_{k}}$ before editing. After editing, the fact chain changes to $\mathcal{P^{*}}=\langle(s_{1}^{L_{1}},r_{1}^{L_{1}},o_{1}^{{L_{1}}*})$ $,(s_{2}^{L_{2}},r_{2}^{L_{2}},o_{2}^{{L_{2}}*}),\dots,(s_{n}^{L_{k}},r_{n}^{L_{k}},o_{n}^{{L_{k}}*})\rangle$ since edits in the first fact can affect the subsequent facts it’s linked to. For answering $Q_{L_{n}}$ after editing, the model has to account for this rippling effect, and provide the final answer as $o_{n}^{{L_{k}}*}$. For this, model has to transfer knowledge of the edited fact and the answer, between languages $L_{1\dots k}$ and $L_{n}$, while correctly reasoning about fact edits via $\mathcal{P^{*}}$.
@@ 3 CroLin-MQuAKE Benchmark
We develop one of the first parallel cross-lingual and multi-hop benchmarks for measuring the knowledge editing capabilities of the existing approaches. A parallel benchmark across languages has the same test examples across all the languages, enabling a direct comparison between them. For this, we use existing datasets measuring the multi-hop model editing in English: MQuAKE-CF and MQuAKE-T released by Zhong et al. (2023), which have counterfactual edits and real-world temporal edits respectively. We translate one fact edit per example in these datasets using Google Translate (Google, ) into 7 languages with diverse writing scripts across medium to high resourcedness - German, Spanish, Chinese, Russian, Hindi, Bengali, Swahili. This results in the benchmark: Cross-Lingual Multi-Hop QnA for Knowledge Editing (CroLin-MQuAKE). It has two datasets, CroLin-MQuAKE-CF and CroLin-MQuAKE-T, each having 8 languages, and 3k and 1.8k parallel examples (same examples in all languages) per language, respectively. The translations are verified by human experts proficient in particular languages and evaluation of BLEU score Papineni et al. (2002) using backtranslation. We find that the translation is highly accurate, since we study medium to high resource languages. See Section A.1 for more details.
@@ 4 Benchmark Analysis on Cross-Lingual Multi-hop Knowledge Editing
Evaluation Metrics. We use multi-hop accuracy proposed by Zhong et al. (2023) which measures the accuracy of the final answer of a multi-hop question. We also adopt hop-wise answering accuracy for checking the correctness of intermediate reasoning steps, as proposed by Gu et al. (2024a).
@@
  • PokeMQA-CL: PokeMQA (Gu et al., 2024a) is similar to MeLLo but consists of a conflict disambiguator for retrieving as well as classifying if a fact is useful to answer a sub-question. Following PokeMQA, we train this disambiguator using BCE loss with negative sampling for retrieving the close edits, given a decomposed sub-question. However, our training dataset now consists of translated version of the training dataset used in PokeMQA. This training set contains all 8 languages (the multilingual setting) or English along with one of the 7 non-English languages (the bilingual setting).
@@
Knowledge editing performance based on retriever training technique. MeLLo-CL retrieves the edited fact from the memory using mContriever and employs an LLM to disambiguate between the generated answer and the retrieved fact and hence ascertains if the generated fact needs any update or not. On the other hand, the current state-of-the-art knowledge editing method in English, PokeMQA-CL, uses a retrieve-then-verify approach, which offloads the knowledge disambiguation to the retriever. This retriever is a light-weight and fine-tuned distilbert-base model trained on a (sub-question,edit) pair dataset using binary cross-entropy loss with negative sampling. It retrieves the closest edits (in fact memory) to a sub-question and scores it on whether the edit answers the question or not (called verification or disambiguation). If it does, then it uses this new knowledge as the answer to the sub-question in the n-th hop step and performs in-context editing. PokeMQA-CL outperforms MeLLo-CL on in the monolingual (English) setting, with a much smaller retriever as shown in Gu et al. (2024a), however, when trained with multilingual data, we find that it significantly under-performs MeLLo-CL in most languages including English as shown in Fig. 2. MeLLo-CL under-performs in Hindi and Bengali – languages with scripts very different from Latin, even though its retriever is trained with 100+ languages.
@@
❶ MeLLo-CL: When edits are made in English, 63.3% of the samples are correct, 29.3% have the contradiction error, 3.6% have Incorrect retrieval, and 3.6% have the incorrect LLM response. For edits made in Hindi, 33.3% of the samples are correct, 60% exhibit an error combination of incorrect retrieval and subsequent contradiction error, where the model first makes an incorrect retrieval and then fails in the contradiction step and 6.6% of erroneous samples are due to the incorrect LLM response. In the CroLin-MQuAKE-CF case when the multilingual edited fact memory contains edits in English and Hindi, MeLLo-CL’s retriever rarely retrieves edits in Hindi, indicating a limitation in its multilingual capabilities. The limitation of MeLLo-CL lies in its retriever-then-contradict mechanism which is up to the LLM.
❷ PokeMQA-CL: When edits are made in English, 53.3% of the samples are correct and 46.3% have the incorrect retrieval error. When edits are made in Hindi, 43.3% are correct, 51% have errors due to the incorrect retrieval and 5.6% are due to the incorrect LLM response. The limitation of PokeMQA-CL lies in its reliance on a bag-of-words model for retrieval. For instance, when presented with the sub-question "Who is the head of state of the USA?", it retrieves the fact “The head of state of Mongolia is Khürelsükh Ukhnaa." This example underscores that PokeMQA-CL prioritizes facts with the highest word overlap, specifically “head of state" indicating a superficial word matching rather than a contextual grasp of the entities involved.
@@ 5 CLeVer-CKE for Knowledge Editing
For overcoming limitations in cross-lingual multi-hop knowledge editing, we design CLeVer-CKE, a cross-lingual and light-weight model editor that seamlessly integrates into any backbone LLM, without changing its parameters. CLeVer-CKE is inspired by memory-based and retrieval-augmented knowledge editing methods (Zhong et al., 2023; Gu et al., 2024a; Mitchell et al., 2022b) for mutlihop question answering. CLeVer-CKE follows the following procedure: Given an input query, it a) decomposes the multi-hop question into multiple sub-questions for getting to the final answer, and for answering each sub-question b) retrieves a relevant fact from the edit memory, c) disambiguates whether the retrieved new knowledge is relevant to answering the sub-question, and d) continues the model generation process based on that. In this work, we primarily aim at showing the importance of having a high-quality retriever for the retrieve-and-verify steps at b) and c) described as follows. See Fig. 3 for an overview.
@@
BCE Loss: Following (Gu et al., 2024a; Mikolov et al., 2013) we add a binary cross-entropy loss in the cross-lingual setting as a baseline loss for training the retriever for retrieving edits in a cross-lingual setting. The negative BCE Loss function takes questions in English and their corresponding edited facts in one of the seven languages as input. We then compute the $L_{2}$ norm between these edits and questions, and sample 20 negatives. The loss function $\mathcal{L}$ is defined similar to Gu et al. (2024a):
@@ 6 Related Works
Knowledge editing methods: Knowledge editing can be broadly classified intro two groups. 1) Parameter-modifying based editing which locates the parameters related to factual knowledge and subsequently modify them (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022a; Meng et al., 2022a, b). These method requires an error-prone analytic step to identify parameters, which might be model-specific and not efficient. 2) Parameter-preserving based editing keeps the model parameters frozen and explicitly stores the fact edits in an external memory, for retrieval and external validation (Zhong et al., 2023; Gu et al., 2024a; Mitchell et al., 2022c; Hartvigsen et al., 2023). Some recent works like that of Hernandez et al. (2023) have also explored a decoding time approach for editing knowledge. Further, knowledge editing is also explored in multimodal settings, such as for text-to-image models (Basu et al., 2023, 2024; Xiong et al., 2024; Gu et al., 2024b).
@@ 8 Limitations
9 Acknowledgements
This work is, in part, supported by NSF (#CNS-2431516). We thank the anonymous reviewers for their constructive feedback on this work. The views, opinions, and/or findings contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Appendix
A.1 Verification of Translated Data in CroLin-MQuAKE
A.1.1 Human Verification of Translation
@@
A.1.2 Verification of Translations via Backtranslation
@@
A.2 Training Details
We employ the training dataset to train the retriever component of the CLeVer-CKE framework, using the same training set as utilized in training PokeMQA-CL Gu et al. (2024a). Subsequently, we translate this dataset into seven other languages and generate hard negatives following the method outlined in Section 5. The training dataset contains 6688 samples along with translations into 8 langugaes and hard-negative pairs for each edit in the dataset, both of which is created by us for training CLeVer-CKE’s retriever. For training the multilingual retriever, we utilize data from all languages, while for training the bilingual retriever, we focus on English and the target language data. To optimize our method’s performance, we conduct hyperparameter tuning on a validation set derived from CroLin-MQuAKE-CF, comprising 50 samples exclusively for this purpose without involvement in inferencing tasks. The hyperparameters used for tuning are mentioned in Table 4. Our experiments are expensive (See Appendix  A.6) and we do not perform experiments on multiple seeds.
A.3 Method Details